Starting on 2011, and continuing until 2022, a company will repay a debt with payments of Rs. 250 000 annually. What equal annual deposits in a fund paying 5% annually beginning on 2001, and continuing until 2022, are required to repay the debt?


1
Expert's answer
2020-07-26T16:18:05-0400

The total debt from 2011-2022 will amount to:

"250,000\times 12= 3,000,000"

Using the future value formulas, the deposit is computed as follows:

The present value of an annuity due will be composed as follows:

"\text{Present value of an annuity due}=\text{Annuity}\times \dfrac{1-(1+r)^{-n}}{r}\times (1+r)"


"3,000,000=\text{Annuity}\times \dfrac{1-(1+0.05)^{-22}}{0.05}\times (1+0.05)"


Therefore, the annuity will be equivalent to:


"\text{Annuity}=\dfrac{3,000,000}{13.82115271}=217,058.6"